OpenStudy (anonymous):

Find the distance between (1, -2) and (-4, 3). A. 5.83 B. 3.16 C. 7.07 D. 5

OpenStudy (anonymous):

Points given : (1, -2) and (-4 , 3) Formula for fimding the distance between two points is sqrt[( y2 -y1)^2 +(x2 -x1)^2] =sqrt[(3 -(-2) )^2 + (-4 -1)^2] = sqrt[( 5)^2 +(-5)^2] =sqrt[( 25+25)] =sqrt(50) =7.07